OpenStudy (anonymous):

Prove that the vector u=(1,2,3,) and v(3,0,-1) are perpendicular.

OpenStudy (anonymous):

show that the dot products equal to 0

OpenStudy (anonymous):

Take the dot product so you get 1*3+2*0+3*(-1)=3+0-3=0 Therefore the vectors are perpendicular.